OpenStudy (anonymous):

Please help... x^2+2x/ 4x^2y - y/14x^2

jimthompson5910 (jim_thompson5910):

this problem is \[\Large \frac{x^2+2x}{4x^2y} - \frac{y}{14x^2}\] correct?

OpenStudy (anonymous):

yes.

jimthompson5910 (jim_thompson5910):

the LCD is 28x^2y do you see how I got this?

OpenStudy (anonymous):

kind of

jimthompson5910 (jim_thompson5910):

the LCM of the denominators will give you the LCD so the LCM of 4x^2y and 14x^2 is 28x^2y Further breakdown: LCM of 4 and 14 is 28 LCM of x^2 and y is x^2y

jimthompson5910 (jim_thompson5910):

so that's how I got the LCD of 28x^2y

OpenStudy (anonymous):

okay i see it now

jimthompson5910 (jim_thompson5910):

the goal is to get each denominator equal to the LCD once you've achieved that goal, you can subtract the two fractions

jimthompson5910 (jim_thompson5910):

so how do we get that first fraction to have a denominator of 28x^2y ?

OpenStudy (anonymous):

so for the first fraction i would multiply it by 7?

jimthompson5910 (jim_thompson5910):

good, multiply top and bottom by 7

OpenStudy (anonymous):

then the second on by 2y?

jimthompson5910 (jim_thompson5910):

\[\Large \frac{x^2+2x}{4x^2y} - \frac{y}{14x^2}\] \[\Large \frac{7(x^2+2x)}{7*4x^2y} - \frac{y}{14x^2}\] \[\Large \frac{7x^2+14x)}{28x^2y} - \frac{y}{14x^2}\]

jimthompson5910 (jim_thompson5910):

very good

jimthompson5910 (jim_thompson5910):

\[\Large \frac{7x^2+14x)}{28x^2y} - \frac{y}{14x^2}\] \[\Large \frac{7x^2+14x)}{28x^2y} - \frac{2y*y}{2y*14x^2}\] \[\Large \frac{7x^2+14x)}{28x^2y} - \frac{2y^2}{28x^2y}\] now you can subtract the numerators and you place the result over the LCD

jimthompson5910 (jim_thompson5910):

that will give you \[\Large \frac{7x^2+14x-2y^2}{28x^2y}\]

OpenStudy (anonymous):

so from there do you simplify the problem to get it to simplest form?

jimthompson5910 (jim_thompson5910):

yes that's only true if you could simplify the numerator, but you can't

jimthompson5910 (jim_thompson5910):

so you just leave it as \[\Large \frac{7x^2+14x-2y^2}{28x^2y}\]

OpenStudy (anonymous):

okay:) thank you very much!

jimthompson5910 (jim_thompson5910):

you're welcome